Find the area of the region inside both the rose curve r=sin⁡〖(2θ)〗 and the circle r=cos⁡θ. Choose the most appropriate graph and intervals to the best of your interest to solve this question
1
Expert's answer
2020-07-09T20:11:40-0400

rose curve r=sin(2"\theta")

circle r=cos"\theta"

"\implies" sin(2"\theta")=cos"\theta"

2sin"\theta" cos"\theta" =cos"\theta"

2sin"\theta"cos"\theta" -cos"\theta" =0

cos"\theta" (2sin"\theta" -1)=0

cos"\theta" =0 , "\theta" = "\pi"/2

2sin"\theta" -1=0

2sin"\theta" =1 , sin"\theta" =1/2 ,"\theta" =("\pi"/6)

now we got the best interval for integrated function is (0,"\pi"/6),("\pi"/6,"\pi"/2)

Area of bounded reason is

A= "\int^{\pi/6}_0" {sin (2"\theta"}2d"\theta" + "\int^{\pi/2}_{\pi/6}" (cos"\theta")2d"\theta"

A= "\int^{\pi/6}_0" 1/2 {1-cos(4"\theta") d"\theta" + "\int^{\pi/2}_{\pi/6}" 1/2 (1+cos(2"\theta")d"\theta"

A=1/2["\theta"-(sin(4"\theta")/4)]"^{\pi/6}_0" + 1/2["\theta"+(sin2"\theta")/2]"^{\pi/2}_{\pi/6}"

A=1/2[{("\pi"/6)-("\sqrt3" )/2/4}-(0-0) + 1/2[{("\pi"/2)+1/2} - {("\pi"/6)+("\sqrt3" )/2/2}

A=[("\pi"/12)-("\sqrt3" )/16] + [("\pi"/4)+(1/4)-("\pi"/12)-("\sqrt3")/8]

A=("\pi"/4)+(1/4)-("\sqrt3" )/16 - ("\sqrt3")/8

A="\pi"/4 + 1/4 -("\sqrt3")/16 - ("\sqrt3")*2/8*2

A="\pi"/4 + 1/4 -("\sqrt3" )/16 - 2("\sqrt3" )/16

A="\pi"/4 + 1/4 + 3("\sqrt3" )/16